OpenStudy (anonymous):

2 tanA=3 tan B Prove that tan(A-B)=sin 2B/5-cos2B

Parth (parthkohli):

\[\tan( A - B) = \dfrac{\tan A - \tan B}{1 + \tan A \tan B}\]Since \(\tan A = \frac{3}{2}\tan B\), you can plug it in.\[= \dfrac{\frac{1}{2}\tan B}{1 + \frac{3}{2}\tan ^2 B} = \dfrac{\tan B}{2 + 3 \tan^2 B}\]

OpenStudy (anonymous):

tanA={3 tan B/ 2 } ------------------(1) Prove that: tan(A-B)=sin 2B/5-cos2B LHS : tan (A-B) = tanA – tanB / 1 +tanA tanB = (3 tanB/2) – tanB / 1 + tanB *3tanB/2 [From(1)] = [(3tanB – 2tanB)/2 ] / [(2 +3 tan2B)/2 ] = { [(tanB)/2 ]* 2 } / [(2 +3 tan2B) ] =(tanB) / (2 + tan2B) =(sinB/cosB) / (2 +{ 3sin2B/cos2B} ) =(sinB/cosB) / { ([2 cos2B +3sin2B]/cos2B ) } =(sinB* cos2B /cosB) / [2 cos2B +3sin2B] =(sinB*cosB) / [2 cos2B +3sin2B] Multiplying the NUMERATOR and DENOMINATOR by “2” ={ 2(sinB*cosB) }/ { 2[2 cos2B +3sin2B]} = {sin2B} /[ 4cos2B +6sin2B] (2sinB*cosB = sin2B) = {sin2B} /[ 4cos2B +6(1- cos2B) ] = {sin2B} /[ 6- 2cos2B) ] = {sin2B} /[ (1+5- 2cos2B) ] = {sin2B} /[ 5+{1- 2cos2B}] = {sin2B} /[ 5 - { 2cos2B-1}] (2cos2B-1 = cos2x) = {sin2B} /[ 5 - { cos22B}] = sin2B / 5 - cos22B LHS = RHS (Hence proved)